40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1667     Entry time: Thu Jun 11 03:15:15 2009
Author: Alberto 
Type: Update 
Category: LSC 
Subject: DD Handoff attempts 

Pete, Alberto,

tonight we worked on the tuning of the double demod phases for the handoff of the short DOFs control signals.

Only MICH can now undergo the handoff. PRC can't make it.

Basically, we tuned the PD6 demod phase and reduced the offset in PD6_I. Then we tuned the relative gain of PD6_I and PD2_I so that the two open loop transfer function of the control loops would match. We tried that in several ways and several times but without success.

I guess we're missing to do/check something.

ELOG V3.1.3-